A company is considering 3 solar panel installations across the city for electric power generation. all analysis is to be done based on rate of return.

Solar Panel 1: Installation fee is $550K. Annual operating cost is $120K first two years and decreases by 5% from year 3. the revenue expected to ear $200K per year from year 2 and increases by 20K yearly. the salvage value is $125K

Solar Panel 2: Installation fee is $830K. Annual operating cost is 120K and increases by 10% annually. the revenue is expeted to be 250K per year and increases $20K yearly. the salvage value is $240K

Solar Panel 3: Installation fee is $830K. Annual operating cost is $120K and increases by 10% annually. the revenue expected to be $300K first two years and increases by $20K from year 3. Additionally there will be a $250K maintenance cost at year 7. the Salvag value is 300K

all solar panels are expected to last 10 years

A) determine the Rate of Return of each project. Show all work including determining number of possible i values and listing all i* values.

B) if MARR is 15% which solar panel option should the company select?
Answer #1

Answer

Option 1

All Amount in 000$ Years 5 Particulars 0 1 2 3 4 6 8 9 10 Installation Cost (A) -550 Expected Revenue | 0 0 220 240 260 280

As we know that IRR be the rate of interest on which present value of cash inflow is equal to present value of cash out flow i.e. NPV is zero.

So let "r" be the required rate of return on which NPV is zero.

Therefore in the given case

NPV @r% = -550-120/(1+r)^1+100/(1+r)^2+126/(1+r)^3+151.70/(1+r)^4+177.12/(1+r)^5+202.26/(1+r)^6+227.15/(1+r)^7+251.79/(1+r)^8+276.20/(1+r)^9+425.39/(1+r)^10

Put r=18%, We get

NPV @18%=9.24

Again Put r=19%, We get

NPV@19% = -22.12

By the method of interpolation

IRR= 18+9.24/(9.24+22.12)*100= 18.29%

Option 2

Statement of computation of Cash Flows

Particulars 0 1 2 3 4 Years 5 6 7 8 9 10 Installation Cost -830.00 Expected Revenue - 250.00270.00 290.00 310.00 330.00 350.0Similarly as calculated in Option 1

Let " i " be the required rate of return on which NPV becomes zero.

So,

NPV @ i% = -830-130/(1+r)^1+138/(1+r)^2+144.80/(1+r)^3+150.28/(1+r)^4+154.31/(1+r)^5+156.74/(1+r)^6+157.41/(1+r)^7+156.15/(1+r)^8+152.77/(1+r)^9+387.50/(1+r)^10

For Reference, D.F.=1/(1+R)^n

YEAR 0 1 2 3 4 5 6 7 8 9 10 NPV
D.F. @14% 1 0.877193 0.769468 0.674972 0.59208 0.519369 0.455587 0.399637 0.350559 0.307508 0.269744
Dis. Cash Flows     -830.00      114.04      106.19         97.74         88.98         80.14         71.41         62.91         54.74         46.98      104.40         -2.48

NPV @14%=- 2.48

YEAR 0 1 2 3 4 5 6 7 8 9 10 NPV
D.F. @13% 1 0.884956 0.783147 0.69305 0.613319 0.54276 0.480319 0.425061 0.37616 0.332885 0.294588
Dis. Cash Flows     -830.00      115.04      108.07      100.35         92.17         83.75         75.28         66.91         58.74         50.85      114.02         35.20

NPV@13%= 35.20

Now Required Rate of IRR= 14-2.48/(35.20-2.48)*1=14-0.075=13.92%

Option 3

Years Particulars 0 1 2 3 4 5 Installation Cost -830.00 Expected Revenue - 300.00 300.00 320.00 340.00 360.00 380.00 400.00 4

Similarly as calculated in above two cases

Let "K" be the required rate of IRR so that NPV of project becomes zero.

NPV @K%= -830+180/(1+K)^1+168/(1+K)^2+174.80/(1+K)^3+180.28/(1+K)^4+184.31/(1+K)^5+186.74/(1+K)^6-62.59/(1+K)^7+186.15/(1+K)^8+182.77/(1+K)^9+477.05/(1+K)^10

YEAR 0 1 2 3 4 5 6 7 8 9 10 NPV
D.F. @16% 1 0.862069 0.743163 0.640658 0.552291 0.476113 0.410442 0.35383 0.305025 0.262953 0.226684
Dis. Cash Flows     -830.00      155.17      124.85      111.99         99.57         87.75         76.65       -22.15         56.78         48.06      108.14         16.81

NPV@16%=16.81

YEAR 0 1 2 3 4 5 6 7 8 9 10 NPV
D.F. @17% 1 0.854701 0.730514 0.624371 0.53365 0.456111 0.389839 0.333195 0.284782 0.243404 0.208037
Dis. Cash Flows     -830.00      153.85      122.73      109.14         96.21         84.06         72.80       -20.85         53.01         44.49         99.24       -15.33

NPV@17%= -15.33

So, Required IRR=16+16.81/(16.81+15.33)=16.52%

Answer (b)

If MARR is 15%, Then option 1 would be selected as it has highest IRR .

Note :- Discounted cash flow= D.F* Cash Flow